A marketing manager travels 1080 miles in a jet and then another 240 miles by car.
If the car ride takes one hour longer than the jet ride takes, and if the rate of the jet is 6 times the rate of the car, find the time the manager travels by jet and find the time the manager travels by car.

Answer by mananth(16949)

let rate of car be x mph
jet rate = 6x
..
distance by jet = 1080
distance by car = 240
...
Car time - jet time = 1
t=d/r
240/x - 1080/6x = 1
LCD= 6x
(1440-1080)/6x= 1
360/6x =1
360 = 6x
/6
x=60 mph
jet rate = 6*60 = 360 mph
...
d car = 240
rate = 60 mph
t = d/r
t=240/60
t=4 hours by car
...
m.ananth@hotmail.ca